# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v4

@@ IV-A3 Active Exploration
For example, Pinto et al.[114] proposed a curious robot that learns visual representations through physical interaction with the environment rather than relying solely on category labels in a dataset. To address the challenge of interactive object perception across robots with varying morphologies, Tatiya et al.[115] innovatively proposed a multi-stage projection framework that transfers implicit knowledge through learned exploratory interactions, enabling robots to effectively recognize object properties without the need to relearn from scratch. Recognizing the challenge of autonomously capturing informative observations, Jayaraman et al. [116] proposed a reinforcement learning method where an agent learns to actively acquire informative visual observations by reducing its uncertainty about unobserved parts of its environment, using recurrent neural networks for the active completion of panoramic scenes and 3D object shapes. NeU-NBV[117] introduced a mapless planning framework that iteratively positions an RGB camera to capture the most informative images of an unknown scene, using a novel uncertainty estimation in image-based neural rendering to guide data collection towards the most uncertain views. Hu et al.[118] developed a robot exploration algorithm that predicts the value of future states using a state value function, combining offline Monte-Carlo training, online Temporal Difference adaptation, and an intrinsic reward function based on sensor information coverage. To address the issue of accidental input in open-world environments, Fan et al.[119] proposed a strategy that treats active recognition as a sequential evidence-gathering process, providing step-by-step uncertainty quantification and reliable prediction under evidence combination theory while effectively characterizing the merit of actions in open-world environments through a specially developed reward function.
@@ IV-B2 One-stage 3D Visual Grounding methods
In Fig. 8 (c), in contrast to two-stage 3D VG methods, one-stage 3D VG methods integrate object detection and feature extraction guided by language queries, making it easier to locate objects relevant to the language.
@@
However, the aforementioned methods either extract sentence-level features that couple all words or focus more on object names in the description, which would lose the word-level information or neglect other attributes. To address these issues, EDA[145] explicitly decoupled the textual attributes in a sentence and conducted dense alignment between such fine-grained language and point cloud objects. Firstly, the long text is decoupled into five semantic components, including main object, auxiliary object, attributes, pronoun, and relationship. Subsequently, the dense alignment is designed to align all object-related decoupled textual semantic components with visual features. To be able to reason human intentions from implicit instructions, ReGround3D [146] designed a visual-centric reasoning module, powered by a Multi-modal Large Language Model, and a 3D grounding module that accurately obtains object locations by revisiting enhanced geometry and fine-grained details from 3D scenes. Additionally, a Chain-of-Grounding mechanism is employed to improve 3D reasoning grounding through interleaved reasoning and grounding steps.
@@ IV-C2 Method
Else MCR-Agent [188] 2023 Multi-Level Model
@@
Memory-Understanding based. Graph-based learning is a very important part of the methods based on understanding and memory of the past. Graph-based learning usually represents the navigation process in the form of a graph, where the information obtained by the embodied agent at each time step is encoded as nodes of the graph. Embodied agent obtains global or partial navigation graph information as a representation of the historical trajectory. In LVERG [170], the authors encode the language information and visual information of each node separately, design a new language and visual entity relationship graph to model the inter-modal relationship between text and vision, and the intra-modal relationship between visual entities. LM-Nav [174] uses a goal-conditioned distance function to infer connections between original observation sets and construct a navigation graph, and extracts landmarks from the instructions through a large language model, uses a visual language model to match them with the nodes of the navigation graph. Although HOP [175] is not based on graph learning, its method is similar to the graph, requiring the model to model time-ordered information at different granularities, thereby achieving a deep understanding of historical trajectories and memories.
@@
The navigation graph discretizes the environment, but concurrently understanding and encoding the environment is also important. FILM [173] uses RGB-D observations and semantic segmentation to gradually build a semantic map from 3D voxels during the navigation. VER [180] quantifies the physical world into structured 3D units through 2D-3D sampling, providing fine-grained geometric details and semantics.
@@
Different learning schemes explore how to utilize historical trajectories and memories better. Through adversarial learning, CMG [171] alternates between imitation learning and exploration encouragement schemes, effectively strengthening the understanding of instructions and historical trajectories, shortening the difference between training and inference. GOAT [179] directly trains unbiased models through Backdoor Adjustment Causal Learning (BACL) and Frontdoor Adjustment Causal Learning (FACL), conducts contrastive learning with vision, navigation history, and their combination to instructions, enabling the agent to make fuller use of information. The enhanced cross-modal matching method proposed by RCM [172] uses goal-oriented external rewards and instruction-oriented internal rewards to perform cross-modal grounding globally and locally and learns from its own historical good decisions through self-supervised imitation learning. FSTT [177] introduces TTA into VLN and optimizes the model in terms of gradients and model parameters at two scales of time steps and tasks, effectively improving model performance.
@@
The specific application of large models in Memory-Understanding based methods is to understand the representation of historical memory and to understand the environment and tasks based on its extensive world knowledge. NaviLLM [176] integrates the historical observation sequence into the embedding space through the visual encoder, inputs the multi-modal information of the fusion encoding into the large language model and fine-tunes it, reaching the state-of-the-art on multiple benchmarks. NaVid [181] makes improvements in the encoding of historical information, achieves different degrees of information retention on historical observations and current observations through different degrees of pooling. DiscussNav [178] assigns large model experts with different abilities to different roles, drives the large models to discuss before navigation actions to complete navigation decisions, and achieves excellent performance in zero-shot VLN.
@@
Future-Prediction Based. Graph-based learning is also widely used in Future-Prediction based methods. BGBL [184] and ETPNav [187] use a similar method to design a waypoint predictor that can predict movable path points in a continuous environment based on the observation of the current navigation graph node. They aim to migrate complex navigation in a continuous environment to node-to-node navigation in a discrete environment, thereby bridging the performance gap from discrete environments to continuous environments.
@@
Improving the understanding and perception of the future environment through environmental encoding is also one of the research directions for predicting and exploring the future. NvEM [183] uses a theme module and a reference module to perform fusion encoding of neighbor views from the global and local perspectives. This is actually an understanding and learning of future observations. HNR [186] uses a large-scale pre-trained hierarchical neural radiation representation model to directly predict the visual representation of the future environment rather than pixel-level images using three-dimensional feature space encoding, and builds a navigable future path tree based on the representation of the future environment. They predict the future environment from different levels, providing effective references for navigation decisions.
@@
Some reinforcement learning methods are also applied to predict and explore future states. LookBY [182] employs reinforcement prediction to enable the prediction module to imitate the world and forecast future states and rewards. This allows the agent to directly map “current observations” and “predictions of future observations” to actions, achieving state-of-the-art performance at the time. The rich world knowledge and zero-shot performance of large models provide many possibilities for Future-Prediction based methods. MiC [185] requires the large language model to directly predict the target and its possible location from the instructions and provides navigation instructions through the description of scene perception. This method requires LLMs to fully exert its ‘imagination’ and build an imagined scene through prompts.
@@
In addition, there are some methods that both learn from the past and for the future. MCR-Agent [188] designs a three-layer action strategy, which requires the model to predict the target from the instructions, predict the pixel-level mask for the target to be interact, and learn from the previous navigation decision; OVLM [189] requires the large language model to predict the corresponding operations and landmark sequences for the instructions. During the navigation process, the visual language map will be continuously updated and maintained, and the operations will be linked to the waypoints on the map.
IV-D Non-Visual Perception
@@
Tactile sensors provides agents with detailed information about object properties such as texture, hardness and temperature. It allows robots to fully grasp the high-precision tasks in hand, making it vital for agents in the physical world. Tactile perception undoubtedly enhances human-computer interaction and holds great promise [190, 191, 192].
@@ IV-D1 Sensor Design
The principle of human tactile is that the skin changes shape when touched, and its abundant nerve cells send electrical signals [193],which also serves as the basis for designing tactile sensors. Tactile sensor design methods can be divided into three categories: non-vision-based, vision-based, and multi-modal. Non-vision-based tactile sensors, primarily using electrical and mechanical principles, chiefly register fundamental, low-dimensional sensory outputs such as force, pressure, vibration, and temperature[194, 195, 196, 197, 198, 199]. One of the notable representatives is BioTac [200] and its simulator[201]. Vision-based tactile sensors are based on optical principles. Using images of the gel’s deformation as tactile information, vision-based tactile sensors such as GelSight[202], Gelslim[203], DIGIT[204], 9DTact[205], TacTip[206], GelTip[207] and AllSight[208] have been used for numerous applications. Simulations like TACTO[209] and Taxim[210] are also popular. Recent work has focused on cost reduction[205] and integration into robotic hands[204, 211, 212]. Multi-modal tactile sensors, inspired by human skin, combine multi-modal information like pressure, proximity, acceleration, and temperature, using flexible materials and modular design.
@@ IV-D2 Datasets
The datasets of non-vision sensors contain electrode values 3D net force vectors and contact location. Therefore, the objects in the datasets are usually force samples and grasping samples. Its tasks are mainly the estimation of force types, force value and grasping details. The datasets mainly collected by BioTac series[200]. As for vision-based sensors, with their high-resolution images of deformation gel, except estimating the force information and sliding, they focus more on texture classification and 3D reconstruction. The objects in the datasets are usually household object, wildlife environments, different materials and grasping items.At the same time, as the image information can be easily aligned and bound with the other modalities (images, language, audio, etc)[16, 213], the perception of touch in embodied agent mainly revolves around visual-based sensors. The datasets revolves around the Geisight sensors, DIGIT sensor and their simulators[202, 204, 209, 205]. We introduce ten commonly used tactile datasets, which are summarized in Table VIII.
TABLE VIII: Comparison of different vision-based tactile datasets.
@@
BioTac Force Estimation[201] 2019 Non-vision Tac,Force - Real BioTac 20k Samples $\surd$
@@
Metal Material
YCB-Slide[219] 2022 Vision Tac,Vision,Other Daily Necessities BOTH TACTO 10 Objects $\surd$
&DIGIT
ObjectFolder Real[220] 2023 Vision Tac,Vision,Audio Household Object Real Gelslim 100 Objects $\surd$
@@
Metal Material
@@ IV-D3 Methods
In these task, bridging the sim-to-real gap is more than important. Reinforcement learning and GAN-based methods have been proposed to address variations in accurate, on-time robotic manipulation tasks.
@@
GAN-based method. ACTNet[249] proposed an unsupervised adversarial domain adaptation method to narrow the domain gap for pixel-level tactile perception tasks. An adaptively correlative attention mechanism was introduced to improve the generator, which is capable of leveraging global information and focusing on salient regions. However, pixel-level domain adaptation lead to error accumulation, degrade performance, and increased structural complexity and training costs. In comparison, STR-Net[250] proposed a feature-level unsupervised framework for tactile images, narrowing the domain gap for feature-level tactile perception tasks. Moreover, some methods focuse on sim-to-real. For example, the Tactile Gym 2.0 [251]. However, due to its complexity and high cost, it is challenging for practical application.
@@
Figure 11: The top gray box displays the scenes an agent observes during exploration. Below are various types of question answering tasks. Except for the task of answering questions based on episodic memory, the agent ceases exploration once it has gathered sufficient information to answer the question.
Traditional Methods. Various traditional approaches have been employed to enhance tactile representation learning. Autoencoder frameworks have been instrumental in developing compact tactile data representations. Polic et al.[252] used a convolutional neural network autoencoder for dimensionality reduction of optical-based tactile sensor images. Gao et al. [253] created a supervised recurrent autoencoder to handle heterogeneous sensor datasets, while Cao et al.[254] created TacMAE used a masked autoencoder for incomplete tactile data. Zhang et al.[255] introduced MAE4GM, a multimodal autoencoder integrating visuo-tactile data. Since tactile acts as a complement to other modes, Joint Training methods are used to fuse multiple modalities. Yuan et al.[256] trained CNNs with modalities included depth, vision, and tactile data. Similarly, Lee et al. [257] used a variational Bayesian approach for modalities like force sensors series and end-effector metrics. For better learning representation, Self-supervised methods like contrastive learning are also a key technique in binding modalities together. Researches differ in contrastive methods. Lin et al. [258] simply paired tactile inputs with multiple visual inputs and Yang et al. [259] employed visuo-tactile contrastive multiview features. Kerr et al. [218] used InfoNCE loss inspired by CLIP and Guzey et al. [260] used BYOL. These traditional methods have established a solid foundation for tactile representation learning.
LLMs&VLMs Methods. Large Language Models (LLM) and Vision-Language Models (VLM) shows the amazing understanding of cross-modal interactions and strong zero-shot performance recently. Recent works from Yang et al.[191], Fu et al. [221] and Yu et al. [261] encoded and aligned tactile data with visual and language modalities by contrastive pretrained method. Then a large language models like LLaMA would be applied,using fine-tune method to fit tasks like tactile description. The advent of LLM and VLM techniques has further advanced the field, enabling more comprehensive and robust cross-modal tactile representations.

### v5

@@ IV-A3 Active Exploration
For example, Pinto et al.[111] proposed a curious robot that learns visual representations through physical interaction with the environment rather than relying solely on category labels in a dataset. To address the challenge of interactive object perception across robots with varying morphologies, Tatiya et al.[112] proposed a multi-stage projection framework that transfers implicit knowledge through learned exploratory interactions, enabling robots to effectively recognize object properties without the need to relearn from scratch. Recognizing the challenge of autonomously capturing informative observations, Jayaraman et al. [113] proposed a reinforcement learning method where an agent learns to actively acquire informative visual observations by reducing its uncertainty about unobserved parts of its environment, using recurrent neural networks for the active completion of panoramic scenes and 3D object shapes. NeU-NBV[114] introduced a mapless planning framework that iteratively positions an RGB camera to capture the most informative images of an unknown scene, using a novel uncertainty estimation in image-based neural rendering to guide data collection towards the most uncertain views. Hu et al.[115] developed a robot exploration algorithm that predicts the value of future states using a state value function, combining offline Monte-Carlo training, online Temporal Difference adaptation, and an intrinsic reward function based on sensor information coverage. To address the issue of accidental input in open-world environments, Fan et al.[116] treated active recognition as a sequential evidence-gathering process, providing step-by-step uncertainty quantification and reliable prediction under evidence combination theory while effectively characterizing the merit of actions in open-world environments through a specially developed reward function.
@@ IV-B2 One-stage 3D Visual Grounding methods
In Fig. 8 (c), in contrast to two-stage 3D VG methods, one-stage 3D VG methods integrate object detection and feature extraction guided by language queries, making it easier to locate objects.
@@
However, these methods either extract sentence-level features that couple all words or focus more on object names in the description, which would lose the word-level information or neglect other attributes. To address these issues, EDA[142] explicitly decoupled the textual attributes in a sentence and conducted dense alignment between fine-grained language and point cloud objects. Firstly, the long text is decoupled into five semantic components, including main object, auxiliary object, attributes, pronoun, and relation. Subsequently, the dense alignment is designed to align all object-related decoupled textual semantic components with visual features. To reason human intentions from implicit instructions, ReGround3D [143] designed a visual-centric reasoning module, powered by a MLMs, and a 3D grounding module that accurately obtains object locations by revisiting enhanced geometry and fine-grained details from 3D scenes. Additionally, a Chain-of-Grounding mechanism is employed to improve 3D reasoning grounding through interleaved reasoning and grounding steps.
@@ IV-C2 Method
Others MCR-Agent [185] 2023 Multi-Level Model
@@
Memory-Understanding based. Graph-based learning is an essential part of the memory-understanding based method. Graph-based learning usually represents the navigation process in the form of a graph, where the information obtained by the embodied agent at each time step is encoded as nodes of the graph. Embodied agent obtains global or partial navigation graph information as a representation of the historical trajectory. LVERG [167] encoded the language information and visual information of each node separately, design a new language and visual entity relationship graph to model the inter-modal relationship between text and vision, and the intra-modal relationship between visual entities. LM-Nav [171] used a goal-conditioned distance function to infer connections between original observation sets and construct a navigation graph, and extracts landmarks from the instructions through a LLM, uses a visual language model to match them with the nodes of the navigation graph. Although HOP [172] is not based on graph learning, its method is similar to the graph, requiring the model to model time-ordered information at different granularities, thereby achieving a deep understanding of historical trajectories and memories.
@@
The navigation graph discretizes the environment, but concurrently understanding and encoding the environment is also important. FILM [170] used RGB-D observations and semantic segmentation to gradually build a semantic map from 3D voxels during the navigation. VER [177] quantified the physical world into structured 3D units through 2D-3D sampling, providing fine-grained geometric details and semantics.
@@
Different learning schemes explore how to utilize historical trajectories and memories better. Through adversarial learning, CMG [168] alternated between imitation learning and exploration encouragement schemes, effectively strengthening the understanding of instructions and historical trajectories, shortening the difference between training and inference. GOAT [176] directly trained unbiased models through Backdoor Adjustment Causal Learning (BACL) and Frontdoor Adjustment Causal Learning (FACL), conducts contrastive learning with vision, navigation history, and their combination to instructions, enabling the agent to make fuller use of information. The enhanced cross-modal matching method proposed by RCM [169] used goal-oriented external rewards and instruction-oriented internal rewards to perform cross-modal grounding globally and locally and learns from its own historical good decisions through self-supervised imitation learning. FSTT [174] introduced TTA into VLN and optimizes the model in terms of gradients and model parameters at two scales of time steps and tasks, effectively improving model performance.
@@
The specific application of large models in Memory-Understanding based methods is to understand the representation of historical memory and to understand the environment and tasks based on its extensive world knowledge. NaviLLM [173] integrated the historical observation sequence into the embedding space through the visual encoder, inputs the multi-modal information of the fusion encoding into the LLM and fine-tunes it, reaching the state-of-the-art on multiple benchmarks. NaVid [178] maked improvements in the encoding of historical information, achieves different degrees of information retention on historical observations and current observations through different degrees of pooling. DiscussNav [175] assigned large model experts with different abilities to different roles, drives the large models to discuss before navigation actions to complete navigation decisions, and achieves excellent performance in zero-shot VLN.
@@
Future-Prediction Based. Graph-based learning is also widely used in Future-Prediction based methods. BGBL [181] and ETPNav [184] used a similar method to design a waypoint predictor that can predict movable path points in a continuous environment based on the observation of the current navigation graph node. They aim to migrate complex navigation in a continuous environment to node-to-node navigation in a discrete environment, thereby bridging the performance gap from discrete environments to continuous environments.
@@
Improving the understanding and perception of the future environment through environmental encoding is also one of the research directions for predicting and exploring the future. NvEM [180] used a theme module and a reference module to perform fusion encoding of neighbor views from the global and local perspectives. This is actually an understanding and learning of future observations. HNR [183] used a large-scale pre-trained hierarchical neural radiation representation model to directly predict the visual representation of the future environment rather than pixel-level images using three-dimensional feature space encoding, and builds a navigable future path tree based on the representation of the future environment. They predict the future environment from different levels, providing effective references for navigation decisions.
@@
Some reinforcement learning methods are also applied to predict and explore future states. LookBY [179] employed reinforcement prediction to enable the prediction module to imitate the world and forecast future states and rewards. This allows the agent to directly map “current observations” and “predictions of future observations” to actions, achieving state-of-the-art performance at the time. The rich world knowledge and zero-shot performance of large models provide many possibilities for Future-Prediction based methods. MiC [182] required the LLM to directly predict the target and its possible location from the instructions and provides navigation instructions through the description of scene perception. This method requires LLMs to fully exert its ‘imagination’ and build an imagined scene through prompts.
@@
In addition, there are some methods that both learn from the past and for the future. MCR-Agent [185] designed a three-layer action strategy, which requires the model to predict the target from the instructions, predict the pixel-level mask for the target to be interact, and learn from the previous navigation decision; OVLM [186] required the LLMs to predict the corresponding operations and landmark sequences for the instructions. During the navigation process, the visual language map will be continuously updated and maintained, and the operations will be linked to the waypoints on the map.
IV-D Non-Visual Perception: Tactile
@@
Tactile sensors provide agents with detailed information such as texture, hardness and temperature. For the same action, the knowledge learned from vision and tactile sensors may be related and complementary, allowing robots to fully grasp the high-precision tasks in hand. Tactile perception is therefore vital for agents in the physical world and undoubtedly enhances human-computer interaction[187, 188, 189].
@@ IV-D1 Sensor Design
The principle of human tactile is that the skin changes shape when touched, and its abundant nerve cells send electrical signals, which also serves as the basis for designing tactile sensors. Tactile sensor design methods can be divided into three categories: non-vision-based, vision-based, and multi-modal. Non-vision-based tactile sensors, primarily using electrical and mechanical principles, chiefly register fundamental, low-dimensional sensory outputs such as force, pressure, vibration, and temperature[190, 191, 192, 193, 194, 195]. One of the notable representatives is BioTac [196] and its simulator[197]. Vision-based tactile sensors are based on optical principles. Using images of the gel’s deformation as tactile information, vision-based tactile sensors such as GelSight[198], Gelslim[199], DIGIT[200], 9DTact[201], TacTip[202], GelTip[203] and AllSight[204] have been used for numerous applications. Simulations like TACTO[205] and Taxim[206] are also popular. Recent work has focused on cost reduction[201] and integration into robotic hands[200, 207, 208]. Multi-modal tactile sensors, inspired by human skin, combine multi-modal information like pressure, proximity, acceleration, and temperature, using flexible materials and modular design.
@@ IV-D2 Datasets
The datasets of non-vision sensors, mainly collected by BioTac series[196], contain electrode values, force vectors and contact location. Since tasks are mainly the estimation of force and grasping details, objects in datasets are usually force and grasping samples. Vision-based sensors, with high-resolution images of deformation gel, focus more on higher estimation, texture recognition and manipulation. The datasets are collected by Geisight sensors, DIGIT sensors and their simulators[198, 200, 205, 201], consisting household object, wildlife environments, different materials and grasping items. Since image information can be easily aligned and bound with the other modalities (images, language, audio, etc)[14, 209], tactile perception in embodied agent mainly revolves around visual-based sensors. We introduce ten main tactile datasets, which are summarized in Table VIII.
TABLE VIII: Comparison of different tactile datasets.
@@
YCB-Slide[215] 2022 Vision Tac, Vision, Other Daily Necessities Both TACTO&DIGIT 10 Objects $\surd$
ObjectFolder Real[216] 2023 Vision Tac, Vision, Audio Household Objects Real Gelslim 100 Objects $\surd$
@@
TaRF[218] 2024 Vision Tac, Vision Full Scene Simulate DIGIT 19.3k -
@@ IV-D3 Methods
In tactile task, bridging the sim-to-real gap is more than important. Reinforcement learning and GAN-based methods have been proposed to address variations in accurate, on-time robotic manipulation tasks.
@@
GAN-based method. ACTNet[245] proposed an unsupervised adversarial domain adaptation method to narrow the domain gap for pixel-level tactile perception tasks. An adaptively correlative attention mechanism was introduced to improve the generator, which is capable of leveraging global information and focusing on salient regions. However, pixel-level domain adaptation lead to error accumulation, degrade performance, and increased structural complexity and training costs. In comparison, STR-Net[246] proposed a feature-level unsupervised framework for tactile images, narrowing the domain gap for feature-level tactile perception tasks. Moreover, some methods focused on sim-to-real. For example, the Tactile Gym 2.0 [247]. However, due to its complexity and high cost, it is challenging for practical application.
@@
Figure 11: The gray box displays the scenes an agent observes during exploration. The other boxes show various types of question answering tasks. Except for the task of answering questions based on episodic memory, the agent ceases exploration once it has gathered sufficient information to answer the question.
Traditional Methods. Various traditional approaches have been employed to enhance tactile representation learning. Autoencoder frameworks have been instrumental in developing compact tactile data representations. Polic et al.[248] used a convolutional neural network autoencoder for dimensionality reduction of optical-based tactile sensor images. Gao et al. [249] created a supervised recurrent autoencoder to handle heterogeneous sensor datasets, while Cao et al.[250] created TacMAE used a masked autoencoder for incomplete tactile data. Zhang et al.[251] introduced MAE4GM, a multimodal autoencoder integrating visuo-tactile data. Since tactile acts as a complement to other modes, Joint Training methods are used to fuse multiple modalities. Yuan et al.[252] trained CNNs with modalities that included depth, vision, and tactile data. Similarly, Lee et al. [253] used a variational Bayesian approach for modalities like force sensors series and end-effector metrics. For better learning representation, Self-supervised methods like contrastive learning are also key techniques in binding modalities together. Researches differ in contrastive methods Lin et al. [254] simply paired tactile inputs with multiple visual inputs and Yang et al. [255] employed visuo-tactile contrastive multiview features. Kerr et al. [214] used InfoNCE loss and Guzey et al. [256] used BYOL. These traditional methods have established a solid foundation for tactile representation learning.
LLMs&VLMs Methods. LLM and VLM have shown an amazing understanding of cross-modal interactions and strong zero-shot performance recently. Recent works from Yang et al.[188], Fu et al. [217] and Yu et al. [257] encoded and aligned tactile data with visual and language modalities by contrastive pretrained method. Then an LLM like LLaMA would be applied, using a fine-tune method to fit tasks like tactile description. The advent of LLM and VLM techniques has further advanced the field, enabling more comprehensive and robust cross-modal tactile representations.